OpenStudy (anonymous):

A cup has a 5-cm diameter bottom. 7-cm diameter top, and has a height of 9-cm. The volume of the can be approximates by cutting it horizontally into sections exactly with a height if 0.5 cm approximating each section with a cylindrical disk. Then summing the volumes of the disks. What is the Volume? Thank you very much for your help!!

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

 There are 18 disks. The bottom one has a lower diameter of 5 cm and an upper diameter of 5 + (7 - 5)/18 cm = 5.1111 cm. The average diameter is (5 + 5.1111) cm / 2 = 5.0555 cm To find the volume of this disk, treat it like a cylinder with diameter = the average diameter = 5.05555 cm and height = 0.5 cm. Then do this another 17 times, with the same height every time, and the diameter increasing (7 - 5)/18 cm = 0.1111 cm each time. Then add all the volumes together.